# Singularity of the axisymmetric stagnation-point-like solution within a cylinder of the 3D Euler incompressible fluid equations

### v1

@@ Abstract
In this paper we investigate analytically the formation of finite-time singularities in the three-dimensional incompressible Euler equations under the specific ansatz of Gibbon, Fokas, and Doering’s vorticity-stretching model (Gibbon et al., 1999), within a bounded cylindrical domain and under axisymmetric conditions. We derive explicit Lagrangian solutions for the vorticity, its stretching rate, fluid pathlines and velocity components, by utilizing constants of motion associated with the system’s field-dependent infinitesimal continuous symmetries. The central finding is that the existence and nature of a finite-time singularity are determined exclusively by the local geometric structure of the initial vortex stretching rate in the vicinity of its global minimum point. The analysis reveals that whether a singularity forms in a finite time depends on how “flat” the initial vortex stretching rate profile is near the minimum point. Specifically, we demonstrate that the flatter the initial stretching rate is near its minimum, the later the potential singularity occurs, with a sufficiently flat profile completely suppressing a finite-time blowup. For stretching-rate profiles that approach the minimum with a power-law behavior, critical thresholds for the power-law exponent are identified, which separate regular solutions from those that develop a finite-time singularity. These thresholds differ depending on whether the singularity occurs at the cylinder’s centre or on a ring away from the centre, with minima at the centre requiring higher “flatness” to avoid a finite-time singularity of the solution. This work provides a rigorous analytical framework that elucidates how the local geometric structure of the initial conditions governs the potential for singularity formation in 3D fluid flows, offering fundamental insights into the interplay between symmetry, initial data, and the development of extreme events in idealised turbulence.
@@ 1 Introduction
Another field playing an important role in this research is the vorticity $\bm{\omega}:=\nabla\times\bm{u}$, which satisfies a dynamical equation that is obtained by taking the curl of the Euler equation (1):
@@
which is the so-called Helmholtz vorticity equation. From its definition as a curl, it is obvious that $\nabla\cdot\bm{\omega}=0$.
One of the most intriguing topics in mathematical fluid dynamics is whether a finite-time singularity emerges in the solution of the 3D Euler fluid equations (1). Many numerical experiments have revealed the existence of a finite-time singularity, see for example Gibbon (2008); Luo & Hou (2014); Choi et al. (2017); Larios et al. (2018); Hou (2023). A milestone of the investigation of the regularity of the solution to the 3D Euler equations is the so-called BKM theorem (Beale et al., 1984) which points out that the convergence of the time integral of the supremum norm (over the spatial domain) of vorticity is a necessary and sufficient condition for 3D Euler fluids to maintain regularity. Subsequently, Constantin et al. (1996) proved that the finite norm of the velocity field along with vorticity direction vectors of class $C^{1}$ constitute a necessary condition to keep 3D Euler flows regular within a finite time. Over the years, considerable research has been conducted on the study of regularity properties and local well-posedness of the 3D Euler fluid equation, which has been greatly summarised by Drivas & Elgindi (2023). Generally speaking, using a more universal model will broaden the applicability of the results, but it also implies that the description of the behavior of a specific Euler solution will be more abstract. By sacrificing some universality, such as introducing some symmetry class or other restrictive class, some special solutions may be expressed in a more explicit way, making it easier for us to understand the various behaviors of the solutions, including the properties of potentially singular behaviors. See the studies with axisymmetry (Elgindi & Jeong, 2020), self-similarity (Mart´ınez-Argüello & Rica, 2024), and octahedral symmetry (Elgindi & Jeong, 2021), for instance.
@@
As early as 1989, Childress, Ierley, Spiegel and Young exhibited blowup in their numerical analysis of two-dimensional stagnation flow Childress et al. (1989). As for the three-dimensional case, equation (3), the first numerical experiment was implemented by Ohkitani & Gibbon (2000), discovering the finite-time blowup for specific initial conditions and finding the approximate inverse proportional relation $\max_{(x,y)\in\Omega_{2}\subseteq\mathbb{R}^{2}}|\gamma(x,y,t)|\propto(T_{*}-t)^{-c}$ with $c\approx 1$, where $T_{*}$ denotes the singularity time and $\Omega_{2}=\mathbb{T}^{2}$, the $2$-torus. Shortly afterwards, an analytic Lagrangian solution for the stretching rate $\gamma(x,y,t)$ was given by Constantin (2000) in the same case $\Omega_{2}=\mathbb{T}^{2}$. In the limiting case $\Omega_{2}=\mathbb{R}^{2}$ of an infinite domain, Gibbon, Moore and Stuart obtained, in the case of zero plane vorticity $\omega$, explicit axisymmetric solutions for $\gamma$ and pathlines that indeed have finite-time singularity (Gibbon et al., 2003). All these solutions find their generalization in the work of Bustamante (2022) which introduces a rich Lie-algebra structure of infinitesimal contemporaneous symmetries, based on which a general Lagrangian solution of stretching rate and vorticity are attained analytically. Also, to model the corrections that would represent a finite-energy solution, Mulungye, Lucas and Bustamante put forward a pressure field modified on the second derivative in the vertical direction while the Lagrangian solutions are still obtained analytically in terms of a modification parameter (Mulungye et al., 2015). Detailed analysis of the existence and asymptotic behaviors of the singular solutions to the 3D Euler fluid equation under this modified pressure field are illustrated, where the asymptotic form of $\gamma$ is certainly derived as $\max_{(x,y)\in\Omega_{2}\subset\mathbb{R}^{2}}|\gamma(x,y,t)|\propto(T_{*}-t)^{-1}$ (Bustamante, 2022; Mulungye et al., 2015, 2016).
@@
Of note is the numerical experiment by Luo & Hou (2014) of an axisymmetric solution of 3D incompressible Euler equation inside a cylinder (with a slip boundary condition), which disclosed a ring-like singularity located at the cylindrical boundary. Note also that a similar ring-like grow-up (in a domain without boundary $\Omega=\mathbb{T}^{2}\times\mathbb{R}$) was observed by Ohkitani and Gibbon in their numerical simulation of Gibbon’s model (Ohkitani & Gibbon, 2000, Initial condition 2). However, as we will explain at the end of section 4, that configuration does not lead to a finite-time singularity because the local profile of the vorticity stretching rate is ‘flat’ enough at the ring. Therefore, the axisymmetric solution of Gibbon et al.’s vorticity-stretching model for a domain with boundary is worthy of an investigation. Moreover, the existence of a finite-time singularity for the Gibbon et al.’s model is known to depend on the initial condition of stretching rate as an integral (Constantin, 2000; Bustamante, 2022). Nonetheless, it is currently unclear which properties of initial stretching rate, local or global, algebraic or geometric, determine this dependence. On the basis of these views, the singularity of 3D axisymmetric Euler solution under the ansatz of Gibbon et al.’s vorticity-stretching model is discussed in this paper. The existence of the singularity in this case is, as we show in the following, only determined by the local properties of the initial stretching rate in the vicinity of its minimum point.
@@
    Spatial Domain: The spatial coordinates are defined as $\bm{X}=(r,\theta,z)\in\Omega_{C}=[0,R]\times[0,2\pi)\times\mathbb{R}$, indicating that the fluid is bounded by a cylindrical region of radius $R$ and extends infinitely along the vertical $z$-axis. For convenience, we adopt cylindrical coordinates throughout the discussion, where $\hat{r}$ denotes the radial direction centred at the cylinder’s axis ($r=0$), $\hat{\theta}$ represents the azimuthal direction oriented counterclockwise around the central axis, and $\hat{z}$ is the vertical direction;
@@
    Axisymmetry: The fluid is axisymmetric across the entire domain, meaning that both the velocity field $\bm{u}$ and the pressure field $p$ are independent of the azimuthal coordinate $\theta$. Mathematically, this implies:
    $\frac{\partial\bm{u}}{\partial\theta}=\bm{0},\quad\frac{\partial p}{\partial\theta}=0.$
    This setup ensures that all physical quantities depend only on the spatial coordinates $r,z,$ and time $t$, simplifying the analysis while retaining the essential features of the problem.
@@
Gibbon, Fokas, and Doering proposed a fluid model—referred to as Gibbon’s vorticity-stretching model (Gibbon et al., 1999)—to describe the dynamics of vortical structures arising from vortex interactions. This model has been shown to admit singular Lagrangian solutions under xy-periodic boundary conditions. In this work, we extend its application to investigate the formation of singularities in solutions of the three-dimensional (3D) incompressible Euler equations under axisymmetric conditions within a cylindrical domain with one lid.
@@
with $T>0$ denoting the maximal time interval over which the solution remains regular. The velocity field exhibits a stagnation plane at $z=0$, with radial and azimuthal components symmetric about this plane, while the vertical component is antisymmetric. Consequently, it suffices to analyze the fluid behavior in the subregion where $z\geq 0$.
@@
For brevity, we refer to this configuration as the ASCII flow, an acronym for Axisymmetric Stagnation-point-like solution inside a Cylinder of the Inviscid Incompressible 3D fluid equations. We believe that such an acronym reflects the unique character of the solution.
@@
$\bm{\omega}(r,z,t)=z\frac{\partial\gamma}{\partial r}(r,t)\hat{\theta}+\omega(r,t)\hat{z}$, (6)
@@
The term “vorticity stretching rate” (stretching rate, for short) for $\gamma$ in relation to the plane vorticity $\omega$ is justified by the following observation. By considering the two-dimensional restriction of the ASCII flow on the plane $z=0$ from equation (5), we can define the three-dimensional vorticity as $\omega\hat{z}$ and substitute it into the Helmholtz vorticity equation (6), resulting in
$\frac{D\omega}{Dt}=\gamma\omega,$
where $\frac{D}{Dt}$ denotes the material derivative. In this context, $\omega$ represents the vorticity of the reduced two-dimensional fluid. This equation reveals that the rate of change of $\omega$ is directly proportional to $\omega$ itself, with proportionality coefficient $\gamma$, which functions analogously to a zeroth-order strain tensor. Therefore, the quantity $\gamma$ in Gibbon’s velocity model (5) is referred to as the stretching rate.
It is evident from equations (8) and (7) that they represent linear ordinary differential equations in $u_{r}$ and $u_{\theta}$, respectively. Solving these yields explicit expressions for the radial and azimuthal velocity components:
@@
The velocity field is the centrepiece of the Eulerian description of a fluid. Combining equation (9) with the velocity form given in equation (5), we deduce that the Eulerian velocity field can be fully expressed in terms of the plane vorticity $\omega$, its stretching rate $\gamma$, and the vertical coordinate $z$. Consequently, the triplet $(\omega,\gamma,z)$ captures all information necessary for describing the fluid dynamics within the Eulerian framework.
Taking the limit $r\to 0$ in equation (9), and assuming $\gamma$ and $\omega$ are regular fields, we find that
$u_{r}(r=0,t)=u_{\theta}(r=0,t)=0,$
which ensures that no spatial singularity occurs at the axis of symmetry. Furthermore, due to the no-flow boundary condition $u_{r}(r=R,t)=0$, and the relation between $u_{r}$ and $\gamma$ in equation (9), we deduce that
$\langle\gamma(\cdot,t)\rangle=0,$
@@
We note that the system of partial differential equations (7), (8), (11) is invariant under the following scale transformation:
@@
for any nonzero real constant $k$ and arbitrary fields $(u_{r},u_{\theta},\omega,\gamma)$. This scale transformation provides a method to generate an infinite family of new solutions from a known solution through simple rescaling. Of course, this transformation may change the boundary conditions at $r=R$, while it respects those at $r=R/k$.
@@ 2.2 Solutions for Plane Vorticity, Stretching Rate, Pathlines, and Velocity Field
The identification of symmetries and corresponding constants of motion is not the focus of this work, as it has been thoroughly explored by Bustamante (2022). Therefore, we do not elaborate on the derivation of infinitesimal symmetries or their use in constructing conserved quantities. Instead, we present three relevant constants of motion that are essential for our subsequent analysis:
@@
Here, the time-dependent function $S(t)$ is obtained via the solution to a second-order ordinary differential equation:
$\ddot{A}(t)+2\left\langle\gamma^{2}\right\rangle A(t)=0,\quad\text{with}\quad A(t):=\dot{S}(t)^{-\frac{1}{2}},$ (15)
where $\langle\cdot\rangle$, defined in equation (10), denotes the volume-averaged integral over the cylindrical domain.
It is straightforward to verify that these expressions satisfy the conservation law given in Equation (13), confirming their validity as constants of motion. Consequently, along particle trajectories, these quantities remain invariant:
$C_{1}(r(r_{0},t),t)=C_{1}(r_{0},0),\quad C_{2}(r(r_{0},t),t)=C_{2}(r_{0},0),\quad C_{3}(r(r_{0},t),z(r_{0},z_{0},t),t)=C_{3}(r_{0},z_{0},0),$
@@
Without loss of generality, we impose the initial conditions:
$S(0)=0,\quad\dot{S}(0)=1,\quad\ddot{S}(0)=0.$ (16)
Under these conditions, the Lagrangian solutions for the stretching rate $\gamma$ and the plane vorticity $\omega$ are obtained as follows:
@@
where $\gamma_{0}$ and $\omega_{0}$ denote the initial values of the stretching rate and plane vorticity, respectively: $\gamma_{0}(r):=\gamma(r,0)$ and $\omega_{0}(r):=\omega(r,0)$.
These solutions explicitly describe the evolution of vorticity and its stretching rate in terms of the function $S(t)$, which governs the temporal dynamics of the system.
@@
In addition to the initial conditions, the quantities $\gamma$, $\omega$, and $z$ are all functions of a single time-dependent variable $S(t)$, which is a strictly increasing function of time $t$. The evolution of $S(t)$ is governed by the following ordinary differential equation:
@@
$\langle f(\cdot;t)\rangle_{0}:=\frac{1}{\pi R^{2}}\int_{0}^{2\pi}\int_{0}^{R}f(r_{0};t)\,r_{0}\,\mathrm{d}r_{0}\,\mathrm{d}\theta_{0}.$ (19)
This average is taken over the Lagrangian coordinates $(r_{0},\theta_{0})$ at the initial time, distinguishing it from the Eulerian average over current spatial coordinates. This ODE provides a means to compute $S(t)$ explicitly, which in turn allows us to determine other physical quantities such as the stretching rate $\gamma$ and vorticity $\omega$.
@@
The derivation of equation (18) will be presented in the proof of Theorem 1 below.

### v2

@@ Abstract
In this paper we investigate analytically the formation of finite time singularities in the three dimensional incompressible Euler equations under the model of Gibbon, Fokas, and Doering for vorticity stretching within a bounded cylindrical domain and under axisymmetric conditions. We derive explicit Lagrangian solutions for the vorticity, its stretching rate, fluid pathlines, and velocity components by exploiting constants of motion associated with the field dependent infinitesimal symmetries of the system. The central finding is that the existence and nature of a finite time singularity are determined exclusively by the local geometric structure of the initial vortex stretching rate near its global minimum. Whether a singularity forms depends on how flat this profile is at the minimum. Flatter profiles delay the blowup and sufficient flatness can suppress it entirely. For power law behavior near the minimum, critical thresholds for the exponent are identified which separate regular solutions from those that develop a finite time singularity. These thresholds differ depending on whether the singularity occurs at the center of the cylinder or on a ring away from the center, with minima at the center requiring higher flatness to avoid blowup. This work provides a rigorous analytical framework that elucidates how the local geometric structure of the initial conditions governs the potential for singularity formation in 3D fluid flows, offering fundamental insights into the interplay between symmetry, initial data, and the development of extreme events in idealized turbulence.
@@ 1 Introduction
The vorticity, defined as $\bm{\omega}:=\nabla\times\bm{u}$, satisfies the Helmholtz vorticity equation obtained by taking the curl of the Euler equations (1):
@@
From its definition as a curl, it is obvious that $\nabla\cdot\bm{\omega}=0$.
One of the most intriguing topics in mathematical fluid dynamics is whether a finite-time singularity emerges in the solution of the 3D Euler fluid equations (1). Many numerical experiments have revealed the existence of a finite-time singularity, see for example Gibbon (2008); Luo and Hou (2014); Choi et al. (2017); Larios et al. (2018); Hou (2023). A milestone of the investigation of the regularity of the solution to the 3D Euler equations is the so-called Beale-Kato-Majda theorem (Beale et al., 1984) which points out that the convergence of the time integral of the supremum norm (over the spatial domain) of vorticity is a necessary and sufficient condition for 3D Euler fluids to maintain regularity. Subsequently, Constantin et al. (1996) proved that the finite norm of the velocity field along with vorticity direction vectors of class $C^{1}$ constitute a necessary condition to keep 3D Euler flows regular within a finite time. Over the years, considerable research has been conducted on the regularity properties and local well-posedness of the 3D Euler equations; see the recent work by Drivas and Elgindi (2023). While general approaches provide broad theoretical insights, they often do not yield detailed information about the structure or dynamics of potential singularities. In contrast, imposing additional structure—such as axisymmetry (Elgindi and Jeong, 2020), self-similarity (Martínez-Argüello and Rica, 2024), or octahedral symmetries (Elgindi and Jeong, 2021)—can significantly simplify the analysis and make it possible to probe singularity formation with greater precision.
@@
As early as 1989, Childress, Ierley, Spiegel and Young exhibited blowup in their numerical analysis of two-dimensional stagnation flow (Childress et al., 1989). As for the three-dimensional case, equation (3), the first numerical experiment was implemented by Ohkitani and Gibbon (2000), discovering the finite-time blowup for specific initial conditions and finding the approximate inverse proportional relation $\max_{(x,y)\in\Omega_{2}\subseteq\mathbb{R}^{2}}|\gamma(x,y,t)|\propto(T_{*}-t)^{-c}$ with $c\approx 1$, where $T_{*}$ denotes the singularity time and $\Omega_{2}=\mathbb{T}^{2}$, the $2$-torus. Shortly afterwards, an analytic Lagrangian solution for the stretching rate $\gamma(x,y,t)$ was given by Constantin (2000) in the same case $\Omega_{2}=\mathbb{T}^{2}$. In the limiting case $\Omega_{2}=\mathbb{R}^{2}$ of an infinite domain, Gibbon, Moore and Stuart obtained, in the case of zero plane vorticity $\omega$, explicit axisymmetric solutions for $\gamma$ and pathlines that indeed have finite-time singularity (Gibbon et al., 2003). All these solutions find their generalization in the work of Bustamante (2022) which introduces a rich Lie-algebra structure of infinitesimal contemporaneous symmetries, based on which a general Lagrangian solution of stretching rate and vorticity are attained analytically. Also, to model the corrections that would represent a finite-energy solution, Mulungye, Lucas and Bustamante put forward a pressure field modified on the second derivative in the vertical direction, modeled in terms of a free parameter, while Lagrangian solutions are still obtained analytically (Mulungye et al., 2015). Detailed analysis of the existence and asymptotic behaviors of the singular solutions to the 3D Euler fluid equation under this modified pressure field are illustrated, where the asymptotic form of $\gamma$ is certainly derived as $\max_{(x,y)\in\Omega_{2}\subset\mathbb{R}^{2}}|\gamma(x,y,t)|\propto(T_{*}-t)^{-1}$ (Bustamante, 2022; Mulungye et al., 2015, 2016).
@@
Of note is the numerical experiment by Luo and Hou (2014) of an axisymmetric solution of 3D incompressible Euler equation inside a cylinder (with a slip boundary condition), which disclosed a ring-like singularity located at the cylindrical boundary. Note also that a similar ring-like grow-up (in a domain without boundary $\Omega=\mathbb{T}^{2}\times\mathbb{R}$) was observed by Ohkitani and Gibbon in their numerical simulation of Gibbon’s model (Ohkitani and Gibbon, 2000, Initial condition 2). However, as we will explain at the end of section 4, that configuration does not lead to a finite-time singularity because the local profile of the vorticity stretching rate is ‘flat’ enough at the ring. Therefore, the axisymmetric solution of Gibbon et al.’s vorticity-stretching model for a domain with boundary is worthy of an investigation. While it is known that finite-time singularity formation in this model depends on the initial stretching rate through an integral condition (Constantin, 2000; Bustamante, 2022), the precise nature of this dependence—whether governed by local or global features, algebraic or geometric structure—has remained unclear. In this paper, we show that finite-time singularity formation in this model is governed entirely by the local power-law exponent of the initial stretching rate at its minimum: blowup occurs below a critical flatness threshold, and global regularity prevails above it.
Although the stagnation-point-like solution of Gibbon et al. (1999) possesses infinite energy, it captures the essential local dynamics of a vortex tube embedded within a larger, finite-energy flow. The singularity criteria derived here are purely local: they depend only on the geometric structure of the initial stretching rate in a neighborhood of its minimum. This implies that our results can serve as a diagnostic tool to assess whether a localized vortex tube in a realistic flow is prone to finite-time breakdown—after which the tube effectively “disappears” from the perspective of this reduced description.
@@
    Spatial Domain: The spatial coordinates are defined as $\bm{X}=(r,\theta,z)\in\Omega_{C}=[0,R]\times[0,2\pi)\times\mathbb{R}$, indicating that the fluid is bounded by a cylindrical region of radius $R$ and extends infinitely along the vertical $z$-axis. For convenience, we adopt cylindrical coordinates throughout the discussion, where $\hat{r}$ denotes the radial direction centered at the cylinder’s axis ($r=0$), $\hat{\theta}$ represents the azimuthal direction oriented counterclockwise around the central axis, and $\hat{z}$ is the vertical direction;
@@
    Axisymmetry: The fluid is axisymmetric across the entire domain, meaning that both the velocity field $\bm{u}$ and the pressure field $p$ are independent of the azimuthal coordinate $\theta$: $\frac{\partial\bm{u}}{\partial\theta}=\bm{0},\quad\frac{\partial p}{\partial\theta}=0$, on $\Omega_{C}$ and at all times $t\in[0,T)$.
@@
Gibbon, Fokas, and Doering proposed a fluid model—referred to as Gibbon’s vorticity-stretching model (Gibbon et al., 1999)—to describe the dynamics of vortical structures arising from vortex interactions. This model has been shown to admit singular Lagrangian solutions under xy-periodic boundary conditions. In this work, we extend that model to investigate the formation of singularities in solutions of the three-dimensional (3D) incompressible Euler equations under axisymmetric conditions within a cylindrical domain with one lid. Hereafter, the resulting flow will be referred to as the cylindrical Gibbon flow.
@@
with $T>0$ denoting the maximal time interval over which the solution remains regular. We impose that $u_{r}(r,t),u_{\theta}(r,t),\gamma(r,t)$ be continuous on $0,R]\times[0,T)$ and piecewise in $C_{r}^{2}C^{1}_{t}\left([0,R]\times[0,T)\right)$, which denotes the space of functions with continuous second-order partial derivatives with respect to $r$ and continuous first-order partial derivatives with respect to $t$. This prepares us to search for strong solutions of the Euler equation (1) and the Helmholtz vorticity equation (2). The velocity field exhibits a stagnation plane at $z=0$, with radial and azimuthal components symmetric about this plane, while the vertical component is antisymmetric. Consequently, it suffices to analyze the fluid behavior in the subregion where $z\geq 0$.
@@
$\bm{\omega}(r,z,t)=-z\frac{\partial\gamma}{\partial r}(r,t)\hat{\theta}+\omega(r,t)\hat{z}$, (6)
@@
The term “vorticity stretching rate” (stretching rate, for short) for $\gamma$ in relation to the plane vorticity $\omega$ is justified by the following observation. By considering the two-dimensional restriction of the cylindrical Gibbon flow on the plane $z=0$ from equation (5), we can define the three-dimensional vorticity as $\omega\hat{z}$ and substitute it into the Helmholtz vorticity equation (6), resulting in
${\left(\frac{D\omega}{Dt}=\right)\frac{\partial\omega}{\partial t}+u_{r}\frac{\partial\omega}{\partial r}=}\gamma\omega,$
where $\frac{D}{Dt}:=\frac{\partial}{\partial t}+u_{r}\frac{\partial}{\partial r}+u_{z}\frac{\partial}{\partial z}$ denotes the material time derivative. In this context, $\omega$ represents the vorticity of the reduced two-dimensional fluid. This equation shows that the material rate of change of $\omega$ is directly proportional to $\omega$ itself, with proportionality coefficient $\gamma$. Hence, the quantity $\gamma$ in Gibbon’s model (5) is referred to as the stretching rate.
It is evident from equations (8) and (7) that, at a given time $t$, they represent linear ordinary differential equations in $u_{r}$ and $u_{\theta}$, respectively. Solving these yields explicit expressions for the radial and azimuthal velocity components:
@@
The velocity field is the centerpiece of the Eulerian description of a fluid, a role that the pathlines play in the Lagrangian description. Combining equation (9) with the velocity form given in equation (5), we deduce that the Eulerian velocity field can be fully expressed in terms of the plane vorticity $\omega$, its stretching rate $\gamma$, and the vertical coordinate $z$. Consequently, the triplet $(\omega,\gamma,z)$ captures all information necessary for describing the fluid dynamics within the Eulerian framework.
From the requirement of physical consistency of axisymmetric flows in cylindrical coordinates, the velocity must be single-valued at the axis $r=0$, which implies
$u_{r}(r=0,t)=u_{\theta}(r=0,t)=0,\quad t\in[0,T)\,$,
which we impose as boundary conditions from the outset and is consistent with equation (9) when taking the limit $r\rightarrow 0$, thanks to the continuity assumptions on $\gamma$ and $\omega$. These conditions eliminate the coordinate singularity at $r=0$ and are standard for physically admissible axisymmetric Euler flows, see (Batchelor, 2000).
Furthermore, due to the no-flow boundary condition $u_{r}(r=R,t)=0$, and the relation between $u_{r}$ and $\gamma$ in equation (9), we deduce that
$\langle\gamma(\cdot,t)\rangle=0,\quad t\in[0,T)\,$,
@@
It is worth remarking that the term $2\langle\gamma^{2}\rangle$ in the evolution equation for $\gamma$ is a nonlocal term that turns the system (8), (11) into a system of partial integro-differential equations. This nonlocal term originates from the pressure field’s second derivative with respect to the $z$ coordinate. Because the vertical velocity is linear in $z$, consistency of the Euler equations in all three directions requires that the combination $\partial_{t}\gamma+u_{r}\partial_{r}\gamma+\gamma^{2}$ be spatially uniform (i.e., independent of $r$). This uniform value is then determined by taking the volume average over the cross-section. Using the fact that $\langle\gamma\rangle=0$—a consequence of the incompressibility condition together with the no-flow boundary condition at r=Rr=R—the advective term averages to $\langle\gamma^{2}\rangle$, leaving precisely $2\langle\gamma^{2}\rangle$. See (Ohkitani and Gibbon, 2000) for a related derivation.
We note that the system of equations (7), (8), (11) is invariant under the following scale transformation:
@@
for any positive constant $k$ and arbitrary fields $(u_{r},u_{\theta},\omega,\gamma)$. Under this transformation, a solution defined on the cylindrical domain $\Omega_{C}=[0,R]\times[0,2\pi)\times\mathbb{R}$ with no-flow boundary condition $u_{r}(R,t)=0$ is mapped to a new solution on the rescaled domain $\Omega_{C}^{(k)}=\left[0,\frac{R}{k}\right]\times[0,2\pi)\times\mathbb{R}$, satisfying the corresponding boundary condition
$u_{r}^{(k)}(R/k,t)=ku_{r}\left(k\cdot\frac{R}{k},k^{2}t\right)=ku_{r}(R,k^{2}t)=0.$
Thus, the scaling symmetry preserves the form of the problem, including the no-flow condition, provided the cylinder radius is rescaled accordingly. This property allows us to generate an infinite family of solutions from a single reference solution by varying $k>0$.
@@ 2.2 Solutions for Plane Vorticity, Stretching Rate, Pathlines, and Velocity Field
In this subsection we provide a summary of useful results obtained by Bustamante (2022), who carried out a systematic search of infinitesimal contemporaneous symmetries and constants of motion for Gibbon-type vortex stretching models. In that work, the time evolution of the system is encoded in a single scalar function $A(t)$, which satisfies the equation
$\ddot{A}(t)+2\left\langle\gamma^{2}\right\rangle A(t)=0,\quad A(0)=1,\quad\dot{A}(0)=0,\quad t\in[0,T),$ (14)
where $\langle\cdot\rangle$ denotes the volume average over the cylindrical domain (see equation (10)). As we will see below, the definition (15) and the nonlinear ordinary differential equation (19) determine $A(t)$ by quadrature. However, it is useful to think of equation (14) as a linear ordinary differential equation for $A(t)$. This equation establishes a one-to-one correspondence between the pressure’s curvature $p_{zz}(t):=2\left\langle\gamma^{2}\right\rangle$ and the function $A(t)$: for any continuous $p_{zz}(t)$, the initial value problem for (14) admits a unique solution by the Picard–Lindelöf theorem, and conversely, any nontrivial solution $A(t)$ defines a unique $p_{zz}(t)$ via $p_{zz}(t)=2\left\langle\gamma^{2}\right\rangle=-{\ddot{A}}/{A}$. Introducing now the function $S(t)$ via the relation and initial conditions
$S(t):=\int_{0}^{t}\frac{1}{A(t^{\prime})^{2}}dt^{\prime}\quad\Rightarrow\quad S(0)=0,\quad\dot{S}(0)=1,\quad\ddot{S}(0)=0\,$, (15)
one can construct several conserved quantities along fluid particle trajectories. Relation (15) is well defined before the singularity time because, as we will show in theorem 1, the function $S$ satisfies the first-order ordinary differential equation (19), which ensures $\dot{S}>0$ for $t\in[0,T)$. For the purpose of our analysis, we adopt three of the constants of motion derived in Bustamante (2022) (see the paragraphs below equations (3.15) and (3.16) there):
@@
These quantities remain invariant along Lagrangian trajectories: $\frac{DC_{j}}{Dt}=0$ for $j=1,2,3$, as can be verified from (11) after tedious but straightforward calculations. Consequently,
$C_{1}(r(r_{0},t),t)=C_{1}(r_{0},0),\quad C_{2}(r(r_{0},t),t)=C_{2}(r_{0},0),\quad C_{3}(r(r_{0},t),z(r_{0},z_{0},t),t)=C_{3}(r_{0},z_{0},0),$ (17)
@@
Remark 2.2.1
Among the conserved quantities derived in Bustamante (2022), the three constants $C_{1},C_{2},C_{3}$ from equation (16) are particularly useful for our analysis: $C_{1}$ and $C_{2}$ allow us to reconstruct the Lagrangian evolution of $\gamma$ and $\omega$, while $C_{3}$ determines the vertical pathline $z(t)$. Although these constants do not directly correspond to classical physical invariants such as energy or momentum, they encode the geometric structure of the flow along particle trajectories and are essential for constructing the Lagrangian solution.
From the invariant relations (16)–(17), the Lagrangian solutions for the stretching rate $\gamma$ and the plane vorticity $\omega$ are obtained directly. Specifically, the constancy of $C_{2}$ along pathlines gives a formula for the stretching rate $\gamma$, and combining such a formula with the constancy of $C_{1}$ gives a formula for the plane vorticity $\omega$. The result is as follows:
@@
where $\gamma_{0}$ and $\omega_{0}$ denote the initial values of the stretching rate and plane vorticity, respectively: $\gamma_{0}(r):=\gamma(r,0)$ and $\omega_{0}(r):=\omega(r,0)$, and we have used $\dot{S}(0)=1,\,\,S(0)=\ddot{S}(0)=0$. These solutions explicitly describe the evolution of vorticity and its stretching rate in terms of the function $S(t)$, which governs the temporal dynamics of the system.
@@
Along the pathlines, the quantities $\gamma$, $\omega$, and $z$ are all functions of the initial conditions and of a single time-dependent variable $S(t)$, which is a strictly increasing function of time $t$. The evolution of $S(t)$ is governed by the following ordinary differential equation (to be shown in theorem 1):
@@
$\langle f(\cdot,t)\rangle_{0}:=\frac{1}{\pi R^{2}}\int_{0}^{2\pi}\int_{0}^{R}f(r_{0},t)\,r_{0}\,\mathrm{d}r_{0}\,\mathrm{d}\theta_{0}.$ (20)
This average is taken over the Lagrangian coordinates $(r_{0},\theta_{0})$ at the initial time, distinguishing it from the Eulerian average (10) over current spatial coordinates. Despite the presence of an integral, equation (19) is an ordinary differential equation for the scalar function $S(t)$, as the integration is performed over the fixed initial data $\gamma_{0}(r_{0})$ and yields a deterministic function of $S(t)$ alone. Since $\gamma_{0}$ is continuous on the compact interval $[0,R]$, the map $S\mapsto\langle(1+\gamma_{0}S)^{-1}\rangle_{0}$ is smooth for $S<-1/\min\gamma_{0}$ (where $1/\min\gamma_{0}>0$ because $\langle\gamma_{0}\rangle_{0}=0$), and hence its square inverse is locally Lipschitz. The Picard–Lindelöf theorem thus guarantees a unique local solution. Equation (19) provides a means to compute $S(t)$ explicitly, which in turn allows us to determine other physical quantities such as the stretching rate $\gamma$ and vorticity $\omega$.
@@
The derivation of equation (19) will be presented in the proof of Theorem 1 below by integrating formula (26) from $r_{0}=0$ to $R$ and applying the no-flow boundary condition $u_{r}(r=R,t)$. This equation is known in a slightly different context since the work by Constantin (2000), where he uses the notation $\tau$ instead of $S$: see equation (19) there. In its modern form, equation (19) was first introduced by Mulungye et al. (2015) (see equation (3.4) there).